# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ VIII Conclusion
    $\mathrm{CRB}(s_{0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}))=\frac{2s_{0}^{2}}{\alpha(g_{0})}\left[\nabla_{{\rm vec}({\bm{\Sigma})}}^{T}S(\bm{\Sigma}_{0})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0})-\left(\frac{\alpha(g_{0})-1}{(m+2)\alpha(g_{0})-m}\right)\right],$ (191)
    $\bm{\Psi}=2\alpha^{-1}(g_{0})s_{0}\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0}),$ (192)
    Remarkably, it can be shown that, for the scale functional $S_{d}(\bm{\Sigma})\triangleq|\bm{\Sigma}|^{1/m}$ the term $\bm{\Psi}$ cancels out, in accordance with Proposition IV.1. In fact, for $S_{d}(\bm{\Sigma})$, we get $\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}}{m}{\rm vec}(\bm{\Sigma}_{0}^{-1})$ from [23, Ch. 8, Th. 1] and $\mathbf{P}_{S_{d}}(\mathbf{V}_{S,0})=\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\bm{\Sigma}_{0}){\rm vec}(\bm{\Sigma}_{0}^{-1})^{T}$. Consequently $(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}}{m}{\rm vec}(\bm{\Sigma}_{0})$ and thus
@@
    $\mathrm{CRB}(s_{d,0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0}))=\frac{4|\bm{\Sigma}|^{2/m}}{m[m(\alpha(g_{0})-1)+2\alpha(g_{0})]}.$ (193)
  9. .9
    Proof of Proposition VI.1. Let us apply Lemma II.2 to the case of RES distributions for which $\mathcal{T}_{3}=\mathcal{T}_{g_{0}}$. From the definition of $\mathcal{T}_{g_{0}}$ in (75) with the related orthogonal projection in (76), Condition (28) of Lemma II.2 can be expressed as:
    $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})|\mathcal{Q}\right\}=\mathbf{0}.$ (194)
    Then, by using the expression of $\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})$ deduced from (14), we have:
    $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{Q}\right\}=\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\mathbf{I}_{\bm{\xi}_{0}}^{-1}E\left\{\mathbf{s}_{\bm{\xi}_{0}}|\mathcal{Q}\right\},$ (195)
    where $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{Q}\right\}$ and $E\left\{\mathbf{s}_{\bm{\xi}_{0}}|\mathcal{Q}\right\}$ the two sub-vector of $E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}$. To conclude the proof, we just need to evaluate $E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}$. To this end, from the expression of $\mathbf{s}_{\bm{\theta}_{0}}$ given in (110) and by noticing that:
    we immediately have that:
    $\begin{split}E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}&=_{d}-2^{-1}(1+m^{-1}\mathcal{Q}\psi_{0}(\mathcal{Q}))\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\\ &=-2^{-1}(1+m^{-1}\mathcal{Q}\psi_{0}(\mathcal{Q}))[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]^{T}\mathrm{vec}(\bm{\Sigma}_{0}^{-1}).\end{split}$ (197)
    Consequently, by substituting (197) in (195) and by noticing that $\mathrm{Pr}\left(\{1+m^{-1}\mathcal{Q}\psi_{0}(\mathcal{Q})=0\}\right)=0$ (since $\mathcal{Q}$ is a continuous random variable), we get (115).
  10. .10
    Proof of Proposition VI.3. The different steps of the proof are based on some reasoning and notations of [42, Sec. VI of supplement material] for C-CES distributions and the general ”low-rank” model (131), that itself takes up the steps of the proof presented in [49].
@@
    ${\bf T}\triangleq 2^{-1}\alpha(g_{0}){\bf I}_{m^{2}}+4^{-1}(\alpha(g_{0})-1)\mathrm{vec}({\bf I}_{m})\mathrm{vec}({\bf I}_{m})^{T},$ (200)
    which can be written, by means of the notation used in [42, Sec; V of supplement material], in the following compact form
@@
    with ${\bf u}_{n}={\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\mathrm{vec}({\bf I}_{m})={\bf T}^{1/2}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})$. It follows from [49, rel. (14)] that
@@
    where ${\bf W}_{0}^{T}{\bf T}{\bf W}_{0}=\frac{\alpha(g_{0})}{2}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0}\otimes{\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0})+\frac{\alpha(g_{0})-1}{4}{\rm vec}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0}){\rm vec}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0})^{T}$. By noticing that ${\bf W}_{0}^{T}{\bf T}{\bf W}_{0}$ is structured in the form ${\bf A}+{\bf a}{\bf a}^{T}$ and applying the inversion matrix lemma with the trick in (180), we get:
@@
    ${\bf B}_{0}\triangleq\left({\bf I}_{m^{2}}+{\bf K}_{m}\right)\left[\frac{1}{\alpha(g_{0})}({\bf H}_{1,0}\otimes{\bf H}_{1,0})-\frac{\alpha(g_{0})-1}{\alpha(g_{0})(2\alpha(g_{0})+(\alpha(g_{0})-1)p}{\rm vec}({\bf H}_{1,0}){\rm vec}({\bf H}_{1,0})^{T}\right]$ (215)

### v3

@@ VIII Conclusion
    $\mathrm{CRB}(s_{0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}))=\frac{2s_{0}^{2}}{\alpha(\bar{g}_{0})}\left[\nabla_{{\rm vec}({\bm{\Sigma})}}^{T}S(\bm{\Sigma}_{0})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0})-\left(\frac{\alpha(\bar{g}_{0})-1}{(m+2)\alpha(\bar{g}_{0})-m}\right)\right],$ (214)
    $\bm{\Psi}=2\alpha^{-1}(\bar{g}_{0})s_{0}\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0}),$ (215)
    Remarkably, it can be shown that, for the scale functional $S_{d}(\bm{\Sigma})\triangleq|\bm{\Sigma}|^{1/m}$ the term $\bm{\Psi}$ cancels out, in accordance with Proposition IV.1. In fact, for $S_{d}(\bm{\Sigma})$, we get $\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}}{m}{\rm vec}(\bm{\Sigma}_{0}^{-1})$ from [25, Ch. 8, Th. 1] and $\mathbf{P}_{S_{d}}(\mathbf{V}_{S,0})=\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\bm{\Sigma}_{0}){\rm vec}(\bm{\Sigma}_{0}^{-1})^{T}$. Consequently $(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}}{m}{\rm vec}(\bm{\Sigma}_{0})$ and thus
@@
    $\mathrm{CRB}(s_{d,0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0}))=\frac{4|\bm{\Sigma}|^{2/m}}{m[m(\alpha(\bar{g}_{0})-1)+2\alpha(\bar{g}_{0})]}.$ (216)
  11. .11
    Proof of Proposition VI.3. The different steps of the proof are based on some reasoning and notations of [52, Sec. VI of supplement material] for C-CES distributions and the general ”low-rank” model (139), that itself takes up the steps of the proof presented in [57].
@@
    ${\bf T}\triangleq 2^{-1}\alpha(\bar{g}_{0}){\bf I}_{m^{2}}+4^{-1}(\alpha(\bar{g}_{0})-1)\mathrm{vec}({\bf I}_{m})\mathrm{vec}({\bf I}_{m})^{T},$ (219)
    which can be written, by means of the notation used in [52, Sec; V of supplement material], in the following compact form
@@
    with ${\bf u}_{n}={\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\mathrm{vec}({\bf I}_{m})={\bf T}^{1/2}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})$. It follows from [57, rel. (14)] that
@@
    where ${\bf W}_{0}^{T}{\bf T}{\bf W}_{0}=\frac{\alpha(\bar{g}_{0})}{2}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0}\otimes{\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0})+\frac{\alpha(\bar{g}_{0})-1}{4}{\rm vec}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0}){\rm vec}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0})^{T}$. By noticing that ${\bf W}_{0}^{T}{\bf T}{\bf W}_{0}$ is structured in the form ${\bf A}+{\bf a}{\bf a}^{T}$ and applying the inversion matrix lemma with the trick in (203), we get:
@@
    ${\bf B}_{0}\triangleq\left({\bf I}_{m^{2}}+{\bf K}_{m}\right)\left[\frac{1}{\alpha(\bar{g}_{0})}({\bf H}_{1,0}\otimes{\bf H}_{1,0})-\frac{\alpha(\bar{g}_{0})-1}{\alpha(\bar{g}_{0})(2\alpha(\bar{g}_{0})+(\alpha(\bar{g}_{0})-1)p}{\rm vec}({\bf H}_{1,0}){\rm vec}({\bf H}_{1,0})^{T}\right]$ (234)